# Make-It-Animatable: An Efficient Framework for Authoring Animation-Ready 3D Characters

### v2

@@ Abstract
3D characters are essential to modern creative industries, but making them animatable often demands extensive manual work in tasks like rigging and skinning. Existing automatic rigging tools face several limitations, including the necessity for manual annotations, rigid skeleton topologies, and limited generalization across diverse shapes and poses. An alternative approach is to generate animatable avatars pre-bound to a rigged template mesh. However, this method often lacks flexibility and is typically limited to realistic human shapes. To address these issues, we present Make-It-Animatable, a novel data-driven method to make any 3D humanoid model ready for character animation in less than one second, regardless of its shapes and poses. Our unified framework generates high-quality blend weights, bones, and pose transformations. By incorporating a particle-based shape autoencoder, our approach supports various 3D representations, including meshes and 3D Gaussian splats. Additionally, we employ a coarse-to-fine representation and a structure-aware modeling strategy to ensure both accuracy and robustness, even for characters with non-standard skeleton structures. We conducted extensive experiments to validate our framework’s effectiveness. Compared to existing methods, our approach demonstrates significant improvements in both quality and speed. The source code will be made publicly available.
@@
11footnotetext: Permission Pending: Some of the images in this paper are temporarily masked due to copyright issues. We will update them as soon as possible.
@@ 1 Introduction
    Mixamo and Anything World only support simple poses like T- or A-pose, while our methods allow for arbitrary poses.
@@
    Meshy and Mixamo need manual annotation of joint positions.
@@ 2.1 Modeling Dynamics in 3D Vision and Graphics
Figure 2: Pipeline of the proposed framework. Given an input 3D character, we produce high-quality blend weights, bones, and pose-to-rest transformation for it, so that any animation is within easy reach. First, we coarsely localize the joints with a pre-trained lite version of this framework, which enables a finer shape representation. Then the shape is encoded into a neural field with a particle-based autoencoder. The decoding process involves spatial and learnable queries for different animation assets. Finally, the structure-aware modeling of bones is proposed to better align the predictions with skeleton topology priors.
@@ 3.1 Preliminaries
Based on this formulation, given a shape representation of a character composed of $N$ particles, all the assets necessary for animating it (named as animation assets) are: 1) $K$ joints represented by bone head and tail positions, $\bm{J}\in\mathbb{R}^{K\times 6}$, which are connected based on any predefined skeleton topology; 2) a 6-DoF transformation, $\bm{P}\in\mathbb{R}^{K\times 6}$, which transforms the input posed joints to a common rest pose; 3) the blend weights $\bm{W}\in\mathbb{R}^{K\times N}$ between any joint-particle pair. The former two are also referred to as bone attributes since they are orthogonal to the input particles.
@@ 4.2 Comparison Results
Figure 7: Results of more cases to demonstrate the advantage of our method. The detailed explanations can be found in the supplementary material.
@@
In Fig. 4, we compare the proposed method with two commercial software, Meshy [meshy] and Tripo [tripo]. Meshy tends to enlarge the influence regions of skinning weights, leading to incorrect deformations of clothes or muscles. It also fails to extract the rest pose of some complex input shapes and animates the characters in a weird way. Tripo does better in blend weights, but often produces inaccurate and unmatched bones. Plus, it also has difficulty determining the correct rest state for unusual input poses. Moreover, both of the two cannot predict finger bones, resulting in coarse hand poses in the running scenario. In contrast, our method produces complete and accurate bones as well as suitable blend weights, so the animations are also natural and fluid.
In Fig. 5, we evaluate the auto-rigging method RigNet [xu2020rignet] with given test meshes. It can be observed that RigNet often generates unnecessary bone structures, e.g., a cycle graph (the old man case) and a star topology (the bear case). These bone errors are also manifested in Tab. 2. Besides, its weights usually fail to maintain spatial smoothness, resulting in fragmented animating results.
@@
We show more results produced by our method, focusing on the details and some tricky cases. Each sub-figure of Fig. 7 proves some advantages of our methods, which is interpreted one by one as follows. (a) Fine-grained control of fingers. (b) Capacity of abnormal shapes. (c) Complex input poses. (d) Efficiency for high polygon models. (e) Support for asymmetric inputs. (f) Adaptation to non-existing bones. (g),(h) Extension to extra bones.
@@ 5 Conclusion and Discussion
Canonical transformation
Although we have normalized the input shapes to align their scales, they still differ in global transformations. Since the poses we want to predict are relative to the fixed origin, the same pose can have huge numerical differences under different coordinate systems. This will lead to a dramatically increased difficulty in pose prediction. Therefore, we move and rotate the entire shape to a canonical position and orientation before sampling. In practice, we choose this transformation based on an empirically selected datum plane (referred to as the hip plane) determined by three joint positions, i.e., the hip (root of the kinematic tree) and two upper thighs, which can be accurately located even in the coarse stage. The canonical transformation is applied so that: 1) the hip is located at the origin; 2) the normal of the hip plane is aligned with the z-axis; 3) the vector from the right to the left thigh is parallel to the x-axis. This process simplifies the input shape’s spatial distribution and eliminates the pose representation’s ambiguity. Furthermore, the chosen canonical transformation ensures a consistent upright and front-facing orientation of the input character, which is a common prerequisite of many auto-rigging methods [xu2020rignet, ma2023tarig, li2021learning]. By integrating the coarse localization, we now automate this preprocessing step, enabling our framework to effectively handle the inputs regardless of their initial spatial configuration (positions, rotations, and scales).
Hierarchical sampling
Some parts of the input character, e.g., the hands, present fine-grained details within small regions, which typically demand additional resources for accurate processing. However, the uniform sampling applied to the input shape, as well as the subsequent farthest point sampling (FPS) algorithm, usually results in sparsely distributed sample points on hands, which are far from enough to describe the geometry of fingers. Theoretically, increasing the number of sampling points $N$ and the downsampling number $M$ can bring a larger representation capacity for the entire geometry including the hand regions, but that will significantly add to the computational overhead. For efficiency, we choose to keep the total sampling number unchanged and instead leverage the coarse prediction of hand joints. Specifically, we replace the uniform sampling with a hierarchical approach that ensures a designated proportion of sample points are distributed on both hands. Since the FPS algorithm in the shape encoder disrupts the non-uniformity of samples, we adapt it to a hierarchical algorithm as well.
@@
For the shape autoencoder, we use a point cloud of size $N=32768$ as the input. All the points are sampled on the surface of the input mesh. In the fine training or the inference stage equipped with the hierarchical sampling, 50% of the points are uniformly sampled and the other 50% are sampled near the hand joints. The hyperparameters of the shape autoencoder are consistent with the original setting of 3DShape2VecSet [zhang20233dshape2vecset], which internally uses $M=512$ latents (each of them is a $512$-D vector) for the neural field. Our framework is trained on 8 NVIDIA A100 GPUs. The learning rate is linearly increased to $41e-$ 4 within the first 1% iterations (warm-up), and then gradually decreased using the cosine decay schedule until reaching the minimum value of $51e-$ 5.
@@
Figure S4: Additional Comparison with generative 3D methods, i.e., Meshy [meshy] and Tripo [tripo]. We feed them the same image as reference and compare the performance based on their generated 3D models respectively. The blend weights of two joints, i.e., Left Shoulder and Right Leg, are visualized. The T-pose models predicted by our method are included as the front-view animating results. Zoom in to better view the details.
@@
S4.1 Additional Comparison Results
More comparison cases
@@
In addition to Figs. 4 and 6, we exhibit more cases of comparison with the baselines, including Meshy [meshy] and Tripo [tripo] (Fig. S4), TADA [liao2024tada] and HumanGaussian [liu2024humangaussian] (Fig. S5). The results can prove the effectiveness, robustness, and generalizability of the proposed framework.
Comparison with more auto-rigging methods
@@
For comparison with existing auto-rigging methods, we include more cases and compare with two more baselines here in Fig. S6 (in addition to Fig. 5), i.e., Neural Blend Shapes [li2021learning] and TARig [ma2023tarig]. Neural Blend Shapes only supports T-pose inputs and has quite limited generalizability to shapes different from the SMPL mesh. TArig’s skeleton predictions are better than Neural Blend Shapes and RigNet [xu2020rignet], but its blend weights still cannot meet the standard of practical application, producing spatial unsmoothness when deforming the meshes. Besides, all three baselines are unable to produce fine-grained hand bones, while our method handles the fingers well.
@@
Furthermore, the commercial software, Mixamo [mixamo] and Anything World [anythingworld], rely much on the symmetry or pose simplicity (e.g., T-pose and A-pose) of the inputs and will raise errors when faced with complex ones. Therefore, we compare them separately on some additional cases here in Fig. S7. It can be observed that Anything World produces significant errors when extracting the bones of the left arm for the tiger (left case). Meanwhile, Mixamo fails to fail to distinguish the left and right sides of the ninja (right case) and produces a mirrored skeleton. Moreover, when faced with a non-rest input character like this ninja, the predicted pose-to-rest transformations of Mixamo and Anything World both suffer from unnatural deformations, while our results remain good.
Qualitative comparison on the ModelsResource dataset [xu2019predicting]
@@
We also evaluate the proposed framework on the bipedal-humanoid subset of the “ModelsResource-RigNetv1” dataset [xu2019predicting]. Fig. S8 shows some cases for qualitative comparison with RigNet [xu2020rignet] and TARig [ma2023tarig]. Both of these two baselines are trained on the aforementioned dataset, but our model has never encountered a similar data distribution during training. Despite this, our method still achieves the best quality in rigging and skinning, demonstrating its strong generalizability.
S4.2 Additional Visualizations
@@
Figure S9: Results of more cases to demonstrate the advantage of our method. (a) Fine-grained control of fingers; (b) Capacity of abnormal shapes; (c) Complex input poses; (d) Efficiency for high polygon models; (e) Support of asymmetric inputs; (f) Adaptation to non-existing bones; (g) & (h): Extension to extra bones (e.g., long ears and tails).
@@
Detailed explanation of Fig. 7
We provide detailed explanation and analysis for each case in Fig. 7 (also included here in Fig. S9 for quick reference) as follows.
@@
(b) Capacity of abnormal shapes.  For those characters with an exaggerated body ratio (e.g., extremely large head, short limbs, etc), our method can adaptively change the bone length to fit the shape. This is intractable for template-based human-generating works.
@@
(g) & (h) Extension to extra bones.  Once fine-tuned with some extra data (Sec. 3.5), our model is capable of producing positions and weights of non-standard bones. Here we show the long tail of a tiger in (g) and the long ears of a bunny in (h), which are all fully animatable.
Geometry-awareness: attention vs. add
@@
Intuitively, simply concatenating or adding the normals to the coordinates when feeding points into the shape encoder can achieve the same goal as our geometry-aware attention. However, we found in practice that such a vanilla injecting strategy often leads to overfitting on the high-quality training mesh normals. The weight prediction depends so much on normals that some regions are influenced by faraway bones just because they have plausible normals, especially when faced with lower-quality meshes (e.g., produced by generative models). We present a typical case in Fig. S10. While injecting normal information via adding resolves the weight corruption problem in the armpit region, it introduces a new issue of incorrect weights on hands. In contrast, the proposed attention mechanism benefits from normal information without any side effects.
Geometry-aware attention score
@@
In addition to the two exemplar cases in Fig. 8, here we visualize the geometry-aware attention score of more characters in Fig. S11. The distribution of high-attention-score points shows patterns with statistical significance. Specifically, regions with possible spatial ambiguity, e.g., inner thigh, armpit, between ears, etc., are affected more by the normal values. This verifies the effectiveness of our geometry-aware attention, as it works exactly in the desired way to adaptively exploit the normal information.
Limitations of SMPL-based rigging
@@
Template-based avatar generation methods [liao2024tada, liu2024humangaussian] benefit a lot from the well-defined SMPL mesh [loper2015smpl, pavlakos2019smplx], enabling the generation of animatable characters with no additional rigging cost. However, as discussed in Secs. 1 and 4.2, one of the unneglectable limitations of these methods is their inability to depart from realistic human shapes. Although TADA [liao2024tada] attempts to address this issue by predicting vertex deformations of the template mesh, it remains constrained by the preset body ratio of SMPL. Fig. S12 demonstrates some practical examples of cartoon characters with exaggerated body shapes that the SMPL model can hardly accommodate. As illustrated in the left part of Fig. S12, the large heads of these characters cannot be fitted by the template meshes, and the interpolated blend weights will definitely be incorrect in the head regions. In the right part of Fig. S12, we exhibit the results generated by two template-based methods using the same text prompt “cartoon brown bear with extremely large head and short legs”. Despite the specific prompting, both methods still produce body ratios resembling those of realistic humans, which limits their applicability in practical scenarios. In contrast, our method offers a promising solution by making bipedal characters of any shape ready for animation.

### v3

@@ Abstract
3D characters are essential to modern creative industries, but making them animatable often demands extensive manual work in tasks like rigging and skinning. Existing automatic rigging tools face several limitations, including the necessity for manual annotations, rigid skeleton topologies, and limited generalization across diverse shapes and poses. An alternative approach is to generate animatable avatars pre-bound to a rigged template mesh. However, this method often lacks flexibility and is typically limited to realistic human shapes. To address these issues, we present Make-It-Animatable, a novel data-driven method to make any 3D humanoid model ready for character animation in less than one second, regardless of its shapes and poses. Our unified framework generates high-quality blend weights, bones, and pose transformations. By incorporating a particle-based shape autoencoder, our approach supports various 3D representations, including meshes and 3D Gaussian splats. Additionally, we employ a coarse-to-fine representation and a structure-aware modeling strategy to ensure both accuracy and robustness, even for characters with non-standard skeleton structures. We conducted extensive experiments to validate our framework’s effectiveness. Compared to existing methods, our approach demonstrates significant improvements in both quality and speed. More demos and code are available at https://jasongzy.github.io/Make-It-Animatable/.
@@ 1 Introduction
    Meshy and Mixamo are semi-automatic rigging methods, requiring manual annotation of joint positions.
@@
    Mixamo and Anything World only support poses close to T- or A-pose, while our methods allow for arbitrary poses.
@@ 2.1 Modeling Dynamics in 3D Vision and Graphics
Figure 2: Pipeline of the proposed framework. Given an input 3D character, we produce high-quality blend weights, bones, and pose-to-rest transformation for it, so that any animation is within easy reach. First, we coarsely localize the joints with a pre-trained lite version of this framework, which helps enable a finer shape representation. Then the shape is encoded into a neural field with a particle-based autoencoder. The decoding process involves spatial and learnable queries for different animation assets. Finally, the structure-aware modeling of bones is proposed to better align the predictions with skeleton topology priors.
@@ 3.1 Preliminaries
Based on this formulation, given a shape representation of a character composed of $N$ particles, all the assets necessary for animating it (named as animation assets) are: 1) $K$ joints represented by bone head and tail positions, $\bm{J}\in\mathbb{R}^{K\times 6}$, which are connected based on any predefined skeleton topology; 2) a 6-DoF transformation, $\bm{P}\in\mathbb{R}^{K\times 6}$, which transforms the input posed joints to a common rest pose; 3) the blend weights $\bm{W}\in\mathbb{R}^{K\times N}$ between any joint-particle pair. The former two are also referred to as bone attributes, which are the abstract representation encompassing all the input particles (instead of being tied to individual particles like the blend weights).
@@ 4.2 Comparison Results
In Fig. 4, we compare the proposed method with two commercial software, Meshy [meshy] and Tripo [tripo]. Meshy tends to enlarge the influence regions of skinning weights, leading to incorrect deformations of limbs. It also fails to extract the rest pose of some complex input shapes and animates the characters in a weird way. Tripo does better in blend weights, but often produces inaccurate and unmatched bones. Plus, it also has difficulty determining the correct rest state for unusual input poses. Moreover, both of the two cannot predict finger bones, resulting in coarse hand poses in the running scenario. In contrast, our method produces complete and accurate bones as well as suitable blend weights, so the animations are also natural and fluid.
In Fig. 5, we evaluate the auto-rigging method RigNet [xu2020rignet] with given test meshes. It can be observed that RigNet often produces unnecessary topological structures, which is also manifested in its bone prediction errors from Tab. 2. Besides, its weights usually fail to maintain spatial smoothness, resulting in fragmented animating results.
@@ 5 Conclusion and Discussion
Canonical transformation.  Although we have normalized the input shapes to align their scales, they still differ in global transformations. Since the poses we want to predict are relative to the fixed origin, the same pose can have huge numerical differences under different coordinate systems. This will lead to a dramatically increased difficulty in pose prediction. Therefore, we move and rotate the entire shape to a canonical position and orientation before sampling. In practice, we choose this transformation based on an empirically selected datum plane (referred to as the hip plane) determined by three joint positions, i.e., the hip (root of the kinematic tree) and two upper thighs, which can be accurately located even in the coarse stage. The canonical transformation is applied so that: 1) the hip is located at the origin; 2) the normal of the hip plane is aligned with the z-axis; 3) the vector from the right to the left thigh is parallel to the x-axis. This process simplifies the input shape’s spatial distribution and eliminates the pose representation’s ambiguity. Furthermore, the chosen canonical transformation ensures a consistent upright and front-facing orientation of the input character, which is a common prerequisite of many auto-rigging methods [xu2020rignet, ma2023tarig, li2021learning]. By integrating the coarse localization, we now automate this preprocessing step, enabling our framework to effectively handle the inputs regardless of their initial spatial configuration (positions, rotations, and scales).
Hierarchical sampling.  Some parts of the input character, e.g., the hands, present fine-grained details within small regions, which typically demand additional resources for accurate processing. However, the uniform sampling applied to the input shape, as well as the subsequent farthest point sampling (FPS) algorithm, usually results in sparsely distributed sample points on hands, which are far from enough to describe the geometry of fingers. Theoretically, increasing the number of sampling points $N$ and the downsampling number $M$ can bring a larger representation capacity for the entire geometry including the hand regions, but that will significantly add to the computational overhead. For efficiency, we choose to keep the total sampling number unchanged and instead leverage the coarse prediction of hand joints. Specifically, we replace the uniform sampling with a hierarchical approach that ensures a designated proportion of sample points are distributed on both hands. Since the FPS algorithm in the shape encoder disrupts the non-uniformity of samples, we adapt it to a hierarchical algorithm as well.
@@
For the shape autoencoder, we use a point cloud of size $N=32768$ as the input. All the points are sampled on the surface of the input mesh. In the fine training or the inference stage equipped with the hierarchical sampling, 50% of the points are uniformly sampled and the other 50% are sampled near the hand joints. The hyperparameters of the shape autoencoder are consistent with the original setting of 3DShape2VecSet [zhang20233dshape2vecset], which internally uses shape latents $\bm{F}\in\mathbb{R}^{M\times C}$ with $M=512$ and $C=512$ for the neural field. Note that in [zhang20233dshape2vecset], $\bm{F}$ can either be learnable embeddings or initialized by applying farthest point sampling (FPS) on the input point cloud. We choose the latter implementation in this work. Our framework is trained on 8 NVIDIA A100 GPUs. The learning rate is linearly increased to $41e-$ 4 within the first 1% iterations (warm-up), and then gradually decreased using the cosine decay schedule until reaching the minimum value of $51e-$ 5.
S2.3 Body Prior Losses
The implementation of the body prior losses (introduced in Sec. 3.4) involves: 1) defining a prior-based bone pair list, and 2) applying $L_{1}$-based loss to their predicted positions or directions. For example, to encourage bone connectivity, we define a list of bone pairs that should be connected based on the predefined topology, select their predicted head-tail positions, and penalize their pair-wise $L_{1}$ distances.
@@
Figure S4: Additional Comparison with generative 3D methods, i.e., Meshy [meshy] and Tripo [tripo]. We feed them the same image as reference and compare the performance based on their generated 3D models respectively. The blend weights of two joints, i.e., Left Shoulder and Right Leg, are visualized. Given that these baselines can only apply preset motions and their rest-pose models cannot be exported, we apply a similar “running” sequence to all the methods for fair comparison. For non-rest cases, the T-pose models predicted by our method are included as the front-view animating results. Zoom in to better view the details.
@@
S4.1 Analysis of Efficiency
We compare the inference speed with inputs of different vertex numbers in Fig. S9. Our method maintains sub-second time cost even with 10,000+ vertices, achieving 1000x and 100x speedups over RigNet [xu2020rignet] and TARig [ma2023tarig] respectively. While TARig accelerates RigNet’s iterative process through feed-forward regression, we further introduce a particle-based shape encoder that bypasses mesh connectivity processing while achieving better performance. This architecture also enables point sub-sampling, making our approach particularly efficient for high-resolution inputs.
Figure S9: Comparison of the inference time cost with increasing number of vertices as the input.
S4.2 Additional Comparison Results
@@
More comparison cases.  In addition to Figs. 4 and 6, we exhibit more cases of comparison with the baselines, including Meshy [meshy] and Tripo [tripo] (Fig. S4), TADA [liao2024tada] and HumanGaussian [liu2024humangaussian] (Fig. S5). The results can prove the effectiveness, robustness, and generalizability of the proposed framework.
@@
Comparison with more auto-rigging methods.  For comparison with existing auto-rigging methods, we include more cases and compare with two more baselines here in Fig. S6 (in addition to Fig. 5), i.e., Neural Blend Shapes [li2021learning] and TARig [ma2023tarig]. Neural Blend Shapes only supports T-pose inputs and has quite limited generalizability to shapes different from the SMPL mesh. TArig’s skeleton predictions are better than Neural Blend Shapes and RigNet [xu2020rignet], but its blend weights still cannot meet the standard of practical application, producing spatial unsmoothness when deforming the meshes. Besides, all three baselines are unable to produce fine-grained hand bones, while our method handles the fingers well.
@@
Furthermore, the commercial software, Mixamo [mixamo] and Anything World [anythingworld], rely much on the symmetry or pose simplicity (e.g., T-pose and A-pose) of the inputs and will raise errors when faced with complex ones. Therefore, we compare them separately on some additional cases here in Fig. S7. It can be observed that Anything World produces significant errors when extracting the bones of the left arm for the character with a tail (left case). Meanwhile, Mixamo fails to fail to distinguish the left and right sides of the ninja (right case) and produces a mirrored skeleton. Moreover, when faced with a non-rest input character like this ninja, the predicted pose-to-rest transformations of Mixamo and Anything World both suffer from unnatural deformations, while our results remain good.
@@
Qualitative comparison on the ModelsResource dataset [xu2019predicting].  We also evaluate the proposed framework on the bipedal-humanoid subset of the “ModelsResource-RigNetv1” dataset [xu2019predicting]. Fig. S8 shows some cases for qualitative comparison with RigNet [xu2020rignet] and TARig [ma2023tarig]. Both of these two baselines are trained on the aforementioned dataset, but our model has never encountered a similar data distribution during training. Despite this, our method still achieves the best quality in rigging and skinning, demonstrating its strong generalizability.
S4.3 Additional Visualizations
@@
Figure S10: Results of more tricky cases to demonstrate the advantage of our method. (a) Fine-grained control of fingers; (b) Capacity of unusual shapes; (c) Complex input poses; (d) Efficiency for high polygon models; (e) Support of asymmetric inputs; (f) Adaptation to non-existing bones; (g) & (h): Extension to extra bones (e.g., long ears and tails).
@@
Tricky cases.  We show more results produced by our method, focusing on the details and some tricky cases. Each sub-figure of Fig. S10 proves some advantages of our methods, which is interpreted one by one as follows.
@@
(b) Capacity of unusual shapes.  For those characters with an exaggerated body ratio (e.g., extremely large head, short limbs, etc), our method can adaptively change the bone length to fit the shape. This is intractable for template-based human-generating works.
@@
(g) & (h) Extension to extra bones.  Once fine-tuned with some extra data (Sec. 3.5), our model is capable of producing positions and weights of non-standard bones. Here we show the long tail of a monkey in (g) and the long ears of a bunny in (h), which are all fully animatable.
@@
Geometry-awareness: attention vs. adding.  Intuitively, simply concatenating or adding the normals to the coordinates when feeding points into the shape encoder can achieve the same goal as our geometry-aware attention. However, we found in practice that such a vanilla injecting strategy often leads to overfitting on the high-quality training mesh normals. The weight prediction depends so much on normals that some regions are influenced by faraway bones just because they have plausible normals, especially when faced with lower-quality meshes (e.g., produced by generative models). We present a typical case in Fig. S11. While injecting normal information via adding resolves the weight corruption problem in the armpit region, it introduces a new issue of incorrect weights on hands. In contrast, the proposed attention mechanism benefits from normal information without any side effects.
@@
Geometry-aware attention score.  In addition to the two exemplar cases in Fig. 7, here we visualize the geometry-aware attention score of more characters in Fig. S12. The distribution of high-attention-score points shows patterns with statistical significance. Specifically, regions with possible spatial ambiguity, e.g., inner thigh, armpit, between ears, etc., are affected more by the normal values. This verifies the effectiveness of our geometry-aware attention, as it works exactly in the desired way to adaptively exploit the normal information.
@@
Limitations of SMPL-based rigging.  Template-based avatar generation methods [liao2024tada, liu2024humangaussian] benefit a lot from the well-defined SMPL mesh [loper2015smpl, pavlakos2019smplx], enabling the generation of animatable characters with no additional rigging cost. However, as discussed in Secs. 1 and 4.2, one of the unneglectable limitations of these methods is their inability to depart from realistic human shapes. Although TADA [liao2024tada] attempts to address this issue by predicting vertex deformations of the template mesh, it remains constrained by the preset body ratio of SMPL. Fig. S13 demonstrates some practical examples of cartoon characters with exaggerated body shapes that the SMPL model can hardly accommodate. As illustrated in the left part of Fig. S13, the large heads of these characters cannot be fitted by the template meshes, and the interpolated blend weights will definitely be incorrect in the head regions. In the right part of Fig. S13, we exhibit the results generated by two template-based methods using the same text prompt “cartoon brown bear with extremely large head and short legs”. Despite the specific prompting, both methods still produce body ratios resembling those of realistic humans, which limits their applicability in practical scenarios. In contrast, our method offers a promising solution by making bipedal characters of any shape ready for animation.
Figure S14: (1) Failure cases of our method. (a) Challenging pose with self-intersection; (b) & (c) Out-of-distribution shapes; (d) Topological defects (unexpected mesh connections). (2) More results produced by our method with non-rest-pose inputs.
More versatile pose inputs.  It is worth noting that many of our evaluations above are based on inputs close to rest poses because the baselines are not specifically designed to handle versatile pose inputs. Specifically, RigNet [xu2020rignet] and TARig [ma2023tarig] cannot transform arbitrary poses into rest poses. Mixamo [mixamo] and Anything World [anythingworld] raise failure without inputs close to A- or T-pose. Meshy [meshy] and Tripo [tripo] perform poorly on non-rest poses (see Fig. 4 and Fig. S4). To show our method’s generalizability of poses, we provide more results with non-rest inputs in Fig. S14 (2).
Failure cases.  Fig. S14 (1) presents some failure cases of our method. Limitations involving challenging poses and out-of-distribution shapes could be resolved by including more targeted training data. As for the topological defects of input meshes, incorporating other methods to refine the 3D geometry might be a promising solution.